Jim has four test scores of 87,90,91, and 95. There is one more test during the semester. He wants to make an A in the class, which means he needs his average to be a 90. All five tests count the same in determining the class grade. What grade does Jim need to make on the fifth test to make an A in the class?
Answer by Theo(13342)

his total score for 4 tests is 363
if he need to get a 90 average, his total score for 5 tests needs to be 5 * 90 = 450
450 - 363 = 87.
if he gets an 87 on the last test, then his total score for all 5 tests will be 450
divide that by 5 and his average will be 90.
solution:
he needs at least an 87 on the last test.

you could also have done this with averages.

the average of the 4 scores is 90.75

the formula would be (4 * 90.75 + x)/5 = 90

you would get (363 + x) / 5 = 90

multiply both sides of this equation by 5 and you get 363 + x = 450

subtract 363 from both sides of this equaitonm to get x = 450 - 363 = 87

the general formula is A = T / N
T is the total score.
N is the number of scores.
A is the average score.

from this, you can also derive T = A * N

from this, you can also derive N = T / A

in this particular case, N is equal to 5 and A is equal to 90, so T is equal to A * N which is equal to 5 * 90 which is equal to 450.

that's the total you're shooting for and you already know that 4 scores is equal to 363, so the difference between the 5 scores and the 4 scores is what you need on the fifth score.